Kiran drove from Tortula to Cactus, a distance of 240 mi. She increased her speed by 8 mi/h for the 392 mi trip from Cactus to Dry Junction. If the total trip took 12 h, what was her speed from Tortula to Cactus?
Answer by stanbon(75887)

Kiran drove from Tortula to Cactus, a distance of 240 mi. She increased her speed by 8 mi/h for the 392 mi trip from Cactus to Dry Junction. If the total trip took 12 h, what was her speed from Tortula to Cactus?
-------
T to C DATA::
dist = 240 miles ; rate = x mph ; time = d/r = 240/x hrs
--------
C to D DATA:
dist = 392 miles ; rate = x+8 mph ; time = d/r = 392/(x+8) hrs
------
Equation:
time + time = 12 hrs
240/x + 392/(x+8) = 12
------
240(x+8) + 392x = 12x(x+8)
------
240x + 1920 = 12x^2 + 96x
------
2x^2 - 144x - 1920 = 0
-----
x^2 - 72x - 960 = 0
x = 83.5 mph (speed from T to C)
------------
Cheers,
Stan H.
------------